# Quantum Vacuum energy as the origin of Gravity

### v2

@@ V Closing Remarks
Having already summarized the unconventional perspective on Gravity developed in this article and its implications in the Abstract and Introduction, we close with some suggestions for further investigations and discuss some open questions. There are currently many open avenues for testing (and falsifying) our main proposals.
$\bullet$   Above we considered implications for the cosmology based on the $\Lambda{\rm CDM}$ model. The energy scales involved are relatively low since $\mu_{0}=k_{B}T_{0}=2.35\times 10^{-4}\,{\rm eV}$ for $T_{0}=2.7$ K. If our proposed energy dependent Newton’s constant ${\cal G}(\mu)$ in equation (43) is correct, then it is in principle possible to confirm it in some bench-top Cavendish like experiments as a function of temperature. Suppose such experiments are carried out in a small range of temperature $\Delta T$ about a fixed temperature $\widetilde{T}$. From (39)
${\cal G}(T)={\cal G}(\widetilde{T})\left(1-\widehat{b}\log(T/\widetilde{T})\right)\approx{\cal G}(\widetilde{T})\left(1-\dfrac{\widehat{b}}{\widetilde{T}}\,\Delta T\right).$ (61)
As discussed above, the effective Newton’s constant decreases with increased temperature. If $\widetilde{T}=300$ K is room temperature, then $\widehat{b}/\widetilde{T}\approx 6\times 10^{-5}$ per degree change $\Delta T$ if $\widehat{b}=0.02$, where this value of $\widehat{b}$ was inferred from the Hubble tension in (53). There are relatively few experiments along these lines, going back as far as 1923 Shaw ; Dmitriev ; Liangzao , wherein weight is measured as a function of temperature for various metals around room temperature. These experiments indeed consistently measure a decrease in the gravitational force as the temperature is increased, as our model predicts. Based on the ideas in this article, this decrease should be independent of the material, in this case a metal, since the Equivalence Principle is built into the formalism, whereas experiments find slightly different relative changes to $G_{N}$ depending on the metal. This could be due to experimental uncertainties as a result of buoyancy, thermal expansion, and other factors which are difficult to take into account. In any case a linear fit of this experimental data in $\Delta T$ according to (61) yield variations of about $10^{-5}$ per degree, which is roughly consistent with our estimate of $\widehat{b}/\widetilde{T}$ for $\widehat{b}\approx 0.02$ and $\widetilde{T}=300$K. [^20] The article Liangzao does not attempt a fit to their data. We made a crude fit of their tabular results to the functional form (61) and found $\widehat{b}/\widetilde{T}$ to be $0.8\times 10^{-5}$ for gold and $1.4\times 10^{-5}$ for nickel, however a closer inspection and interpretation of this data beyond our crude fit is warranted. At lower reference temperatures $\widetilde{T}$, the effect is larger and this motivates carrying out such Cavendish type experiments at lower temperatures.
@@
$\bullet$   If the varying $G_{N}$ proposed in this article is correct, it is testable by other kinds of measurements. Let us speculate on some possible directions or topics for further study. We already mentioned above that in principle this varying $G_{N}$ can be measured in bench-top Cavendish experiments as a function of temperature, and that the very few existing experiments carried out at room temperature are roughly consistent with the parameters we inferred from the Hubble tension. For cosmology, our formula (39) for the scale dependent Newton’s constant ${\cal G}(\mu)$ implies that the effective Newton’s constant is effectively time dependent if one ties $\mu$ to the time-dependent scale factor as in (42). There are known experimental bounds on the time variation of $G_{N}$, usually reported as $\dot{G}_{N}/G_{N}$. Based on (43) one has
$\dfrac{\dot{G}_{N}}{G_{N}}=H\,\widehat{b}$, (62)
@@
where as above $H=\dot{a}/a$. The strongest experimental constraints come from pulsar timing Taylor which constrains $\dot{G}_{N}/G_{N}<2\times 10^{-12}/{\rm yr}$. LIGO provides weaker constraints $10^{-9}/{\rm yr}$ LIGO . These pulsars exist at relatively low redshift $z$. Thus taking $H=H_{0}$ in the above equation (62), one finds $\dot{G}_{N}/G_{N}\approx 7\times 10^{-11}\times\widehat{b}\,/{\rm yr}$. For $\widehat{b}\approx 0.02$, as we proposed based on Hubble tension data, $\dot{G}_{N}/G_{N}\approx 1.4\times 10^{-12}\,/{\rm yr}$ which is very close to the bound based on pulsars. There are many other testable predictions in cosmology in addition to the Hubble tension and those mentioned above. A varying $G_{N}$ may shift CMB acoustic peaks due to modified expansion at $z\sim 1100$, which is testable with Planck or the Simons Observatory data. The minimal scale factor $a_{\rm min}\sim 10^{-22}$ in eq. (46) could produce a primordial gravitational wave spectrum detectable by LIGO or eventually the space based interferometer LISA. A weaker $G_{N}$ at higher redshift $z$ could delay stellar collapse, favoring primordial black holes. In fact, based on (32), for a weaker $G_{N}$, black holes are smaller for a given mass. Very recently, supermassive black holes, with mass on the order of 40 million solar masses have been discovered using the James Webb Space Telescope JWST which were formed at an earlier time than expected, and this may be an interesting topic for further study in the context of a varying $G_{N}$.
$\bullet$   As stated in the Introduction, although the formula (3) is well-motivated by the work ALrhovac1 ; ALrhovac2 , the QFT underlying this formula for $\rho_{\rm vac}$} has not been precisely identified, in particular the zeron with mass $m_{\rm z}$ has not yet been identified as a known particle, though Majorana fermions are viable candidates. This underlying QFT is incomplete in the UV since it relies on the RG beta function (35) for a marginally irrelevant coupling, where the incompleteness is manifested as a Landau pole at very high energies. However we showed that the time evolution of the scale factor $a(t)$ smoothly traverses this pole through a swing rather than a bang. The UV completeness issue here is certainly more tractable than the UV incompleteness of quantum gravity, since it is a UV issue of QFT in flat Minkowski space where such problems are better understood. Landau poles are commonplace in elementary particle physics and condensed matter physics, since they are generic to marginally irrelevant perturbations. See for instance LandauPole1 ; LandauPole2 ; LandauPole3 and references therein. Although QED itself is not relevant to this article, let us mention that it also is not UV complete due to a Landau pole, however in this context it occurs at much higher energies than our proposed $z_{\rm max}$, and is thus commonly viewed as a purely academic issue, and is furthermore complicated by chiral symmetry breaking. (See footnote 12).

### v3

@@ VIII Closing Remarks
Having already summarized the unconventional perspective on Gravity developed in this article and its implications in the Introduction, we close with some suggestions for further investigations and discuss some open questions. There are currently many open avenues for testing (and falsifying) our main proposals, in addition to those considered in the last section.
@@
$\bullet$   The minimal scale factor $a_{\rm min}\sim 10^{-22}$ in eq. (58) could produce a primordial gravitational wave spectrum detectable by LIGO or eventually the space based interferometer LISA. A weaker $G_{N}$ at higher redshift $z$ could delay stellar collapse, favoring primordial black holes. In fact, based on (30), for a weaker $G_{N}$, black holes are smaller for a given mass. Very recently, supermassive black holes, with mass on the order of 40 million solar masses have been discovered using the James Webb Space Telescope JWST , which were formed at an earlier time and smaller than expected, and this may be an interesting topic for further study in the context of the varying $G_{N}$ that we specifically proposed.
$\bullet$   As stated in the Introduction, although the formula (2) is well-motivated by the works ALrhovac1 ; ALrhovac2 , we left aside the issue of identifying the particle with mass $m_{\rm z}$ underlying this formula for $\rho_{\rm vac}$} since we were able to make progress without doing so, and the particle physics involved is beyond the original scope of this article. Nevertheless, we can constrain some of its properties. We were able to constrain its mass from the observed value of $\rho_{\rm vac}$} in (5), and pointed out that this is consistent with Majorana neutrinos as viable candidates, but were not able to make a strong case for this without a more complete theory of the origin of neutrino masses. Based on our comparison of coupling constants with QED, it could be that the particle with mass $m_{\rm z}$ is coupled to a hidden U(1) gauge theory where $\mathfrak{g}$ is a marginally irrelevant fine-structure constant. We assumed that our formula for $\rho_{\rm vac}$} in (2) is valid up to $z_{\rm max}$, which would seem to imply that such a particle does not obtain its mass $m_{\rm z}$ from the Higgs mechanism since electro-weak symmetry breaking occurs at a significantly lower scale of about $160$ GeV, which corresponds to $z\approx 6\times 10^{14}$. In our current understanding of the Standard Model of particle physics, neutrino masses do not arise from the Higgs mechanism since the Standard Model does not have right-handed neutrinos to pair up with the known left-handed neutrinos to provide a mass. Based on this, a massive Majorana neutrino is the most promising candidate for the $m_{\rm z}$-particle, however other possibilities where this is an entirely new particle are certainly not ruled out. This underlying QFT involving $m_{\rm z}$ is incomplete in the UV since it relies on the RG beta function (47) for a marginally irrelevant coupling, where the incompleteness is manifested as a Landau pole at very high energies. However we showed that the time evolution of the scale factor $a(t)$ smoothly traverses this Landau pole through a swing rather than a bang. The UV completeness issue here is certainly more tractable than the UV incompleteness of quantum gravity, since it is a UV issue of QFT in flat Minkowski space where such problems are much better understood.
$\bullet$   In Section VIID we described some not so well-known small scale bench-top experiments on the temperature variation of Newton’s constant which provided some positive support for our model. It is potentially exciting that our new parameter $\widehat{b}$ for cosmology can in principle be measured by these completely different kinds of experiments, and this justifies reproducing these results with more modern experimental techniques.
IX Acknowledgements
We wish to thank Maria Dainotti, Giovanni Montani, and Biagio de Simone of the NAOJ collaboration for pointing out their recent articles and discussions of their results shortly after the first version of this article appeared. We also benefitted from subsequent discussions with Seamus Davis and Eanna Flanagan.